# Hard rigid rods on Husimi trees

### v1

@@ I Introduction
Particle systems with excluded volume interactions only have been studied for much time. Besides a disordered phase, they may also exhibit ordered phases and thus undergo entropy-driven transitions between them. Many different shapes have been considered for the particles, such as hard spheres, squares, triangles, tetraminoes, and so on cmc05 ; ps88 ; fal07 ; b80 ; vn99 ; bsg09 ; m05 . Some systems composed by binary binary and ternary mixtures ternary of hard particles have also been considered in the literature. In this paper, we investigate hard rigid rods, which are frequently associated to liquid crystals vl92 . The study of this particular system started with a seminal paper by Onsager o49 , where he concluded that, at sufficiently high densities and large aspect ratios, the rods would undergo a discontinuous phase transition between an isotropic and a nematic phase. While Onsager considered the rods in continuous space, Flory f56 placed them on a lattice, but still allowing their orientations to be continuous. He solved the problem in a mean-field approximation, finding also a transition between isotropic and nematic phases. Zwanzig z63 studied a system of hard rods in continuum space, but with discrete orientations, finding, as Onsager, a discontinuous isotropic-nematic phase transition.
@@
Here, we are interested in rigid rods placed on lattices, formed by $k$ consecutive monomers aligned in one lattice direction, so that between consecutive monomers of a rod there is a lattice edge. We will call these rods $k$-mers. For the particular case of $k=2$ (dimers) on the square lattice, the entropy of the system in the limit of full coverage has been exactly determined, using both pfaffians k61 ; tf61 and transfer matrices l67 ; wp21 , being $s_{2}^{(sq)}=G/\pi=0.29156\ldots$ the entropy per site, where $G$ is Catalan’s constant. Fully-packed dimers on the triangular lattice have also a long history (see, e.g., Refs. Nagle66 ; Phares85 ; Kenyon00 ; Fendley02 ; Wu06 ), where $s_{2}^{(tri)}=0.42859\ldots$ Fendley02 ; Wu06 . The exact determination of the entropy for dimers in this limit is also possible for other two-dimensional lattices Wu06 , as well as in higher dimensions dc08 . For larger values of $k$, as far as we known, there are no exact results available for the full lattice entropy of $k$-mers placed on regular lattices. However, rather precise estimates have been obtained using transfer matrix methods for trimers gdj07 and other $k$-mers with $2\leq k\leq 10$ rsd23 on the square lattice, which agree with the outcomes from numerical simulations p21 . Semi-analytical solutions of fully-packed rods on generalized Husimi lattices — built with square lattice clusters of effective lateral size $L$ — also provide results in fair agreement with those for the square lattice, when extrapolated to the $L\rightarrow\infty$ limit rso22 . To the best of our knowledgement, there are no analogous results for the triangular lattice reported in the literature.
@@
Turning our attention to the more general case where vacancies may exist between rods placed on two-dimensional lattices, we start mentioning that Monte Carlo simulations demonstrate that no nematic order exists for $k<k_{min}=7$, whereas a continuous isotropic-nematic transition is found for $k\geq 7$ gd07 . Its universality class depends on the number of possible orientations of the rods; and may be Ising mf08 ; mf08p ; fv09 or, more generally, Potts mf08 ; mf08p ; mf08pp . The existence of the nematic phase at intermediate densities was also rigorously established for large $k$’s ds13 . Interestingly, a second transition — from the nematic to a high-density disordered phase — is found at high rods’ densities gd07 , but its nature is not so well known. Although usual Monte Carlo algorithms using local evaporation and deposition moves are very inefficient in the high density region, a new procedure introduced some time ago (simultaneously updating all sites of a strip using transfer matrix procedures) reduced this difficulty k12 ; k13 , indicating a non-Ising transition for $k=7$. Nevertheless, recent results suggest this transition to be discontinuous sdr22 .
@@
The thermodynamic behavior of $k$-mers was also studied on the Bethe lattice (BL) and Bethe-like lattices drs11 . There, a continuous isotropic-nematic transition was found for the four-coordinated lattice ($q=4$) already for $k\geq k_{min}^{(BL)}=4$; in contrast with $k_{min}=7$ obtained for the square lattice gd07 . For BLs with coordination $q\geq 6$, the isotropic-nematic transition is discontinuous and appears, again, for $k\geq k_{min}^{(BL)}=4$; which differs from the continuous transition found (for $k\geq 7$) in the triangular lattice mf08p . Furthermore, the transition from the nematic to the high density isotropic phase is missing in the BL for all rod size $k$ and coordination $q$. We recall that the exact (or numerically exact) solution of a given model on the BL corresponds to the result of the so-called mean-field Bethe approximation on a regular lattice with the same coordination number b82 . So, the differences above lead us to inquire whether the rods’ behavior can be better captured by more elaborated mean-field approaches. In order to address this, in this paper, we study the thermodynamic properties of $k$-mers, for general densities, on the four-coordinated Husimi lattice built with squares and on the six-coordinated Husimi lattice built with triangles. Since loops are present in these Husimi cacti, they are expected to provide better mean-field approximations for the models on the square and triangular lattices than the simpler BL solution h50 ; t76 . This was indeed the case in the full lattice limit analyzed in Ref. rso22 , as already mentioned above. As demonstrated in what follows, the scenario found here is not so different to that just described for the BL; even though, in most cases, we obtain critical parameters closer to those estimated in the literature for the regular lattices, provided by techniques such as Monte Carlo simulations or transfer matrix calculations.
@@
The rest of the paper is organized as follows. In Sec. II the Husimi lattices are defined and the model is solved on the tree built with squares; while the solution on the triangular Husimi cactus is presented in the Appendix. The thermodynamic behavior of rods placed on the square and triangular Husimi lattices are shown and discussed in sections III and IV, respectively. Our final discussions may be found in Sec. V.
II Solution of the model on the Husimi lattice built with squares
The Husimi tree with coordination number $q=4$, as consider here, is just a treelike structure where successive generations of squares are added to a central square, as is shown in Fig 1(a). A similar tree, with coordination $q=6$, can be built up by joining generations of triangles to a central triangle, as done in Fig. 1(b). The solution of a given model in the core of these Husimi trees, in the thermodynamic limit, is called a Husimi lattice (HL) solution of this model in the literature t76 . The square HL (triangular HL) provides a mean-field approximation for the model’s behavior on the square (triangular) lattice.
@@ II Solution of the model on the Husimi lattice built with squares
Figure 1: Husimi trees built with (a) squares and (b) triangles, by starting with a central polygon and adding two generations of them at the surface sites of the previous generations. Following this procedure, such trees always have an elementary polygon [square in (a) and triangle in (b)] at their centers. The four-coordinated Cayley tree obtained by connecting the squares’ centers in (a) is represented by dashed lines.
@@
The solution of the rigid rods model on the square HL will be presented here, while the one for the triangular case will be discussed in appendix A. The model we are analyzing is a generalization of the full lattice limit, which was recently considered by some of us on the square HL rso22 , but now empty sites are allowed and activities $z_{x}=\exp[\mu_{x}/(k_{B}T)]$ and $z_{y}=\exp[\mu_{y}/(k_{B}T)]$ are associated to each site occupied by a monomer in a rod aligned in the $x$ and $y$ direction, respectively. The quantities $\mu_{x}$ and $\mu_{y}$ are the chemical potentials of a monomer, which is part of a $k$-mer, so that we are in the grand-canonical ensemble. Moreover, $k_{B}$ is Boltzmann’s constant and $T$ is the temperature. Although we will restrict ourselves to the thermodynamic properties of the system in the case $z_{x}=z_{y}=z$, for latter convenience, the solution will be presented considering different activities. As expected, in all expressions below the full lattice limit results are recovered in the $z\to\infty$ limit.
@@
As usual in solving models on such lattices, we start obtaining recursion relations for the grand-canonical partial partition functions (ppf’s) of rooted sub-trees for fixed configurations of their root site. In other words, we consider the operation of connecting three sub-trees to three corners of the root square of the sub-tree with an additional generation of squares. We will label the root configuration of the sub-tree according to the number of monomers already attached to the rod reaching it from above. If there is no such rod, this label will be equal to zero. If there is a reaching rod, the label will be the number of monomers already included in the rod with a second component indicating the direction of the rod, $x$ or $y$. Therefore, there are $12k-$ 1 ppf’s, whose first labels are in the range $[0,...,k-1]$. In Fig. 2 the possible sub-tree root configurations are shown.
Figure 2: Possible root configurations of a sub-tree. From the left to the right: no rod reaching on the root site from above ($g_{0}$), a rod in the $x$-direction arriving at the root site ($g_{(i,x)},\,i=1,2,...,k-1$), and a rod in the $y$-direction reaching on the root site ($g_{(i,y)},i=1,2,...,k-1$).
@@
We now proceed obtaining the recursion relations for the ppf’s, by considering the operation of attaching three sub-trees to a new root square. The possible configurations of the new root square for the case of no rod arriving at the new root site are illustrated in Fig. 3.
@@
Figure 3: Contributions to the recursion relation for $g^{\prime}_{0}$. The thick red lines indicate edges of a $k$-mer placed on the new root square.
@@
and the other recursion relations, for ppf’s of sub-trees with rods arriving at the root site, may be obtained using similar combinatorial arguments. The results are:
@@
for $i=1,2,...,k-1$. We notice that, if we write the recursion relations for the ppf’s as $g^{\prime}_{(i,x)}=g_{0}^{3}f_{(i,x)}$ and $g^{\prime}_{(i,y)}=g_{0}^{3}f_{(i,y)}$, the functions $f_{i}$ depend only of the activities and the ratios defined above. It is easy to obtain the $2(k-1)$ recursion relations for the ratios. They are:
@@
The grand-canonical partition function of the system may be obtained considering the operation of connecting four sub-trees with $N$ generations to the central square of the Husimi tree. A convenient alternative is to connect a sub-tree with $N+1$ generations to another one with $N$ generations, so that they share their root sites. This common root site may be empty or occupied by a monomer of a horizontal or vertical rod, thus:
@@
with
@@
and it is implicit that $g_{0}$ corresponds to the ppf of a sub-tree with an empty root site and $N$ generations. The average number of $k$-mers in the horizontal and vertical directions reaching the site of the central square of the tree where the two sub-trees meet are, respectively:
@@
The free energy of the system on the whole Husimi tree, in the thermodynamic limit, is dominated by the surface. The bulk grand-canonical free energy per site, limited to the central region, discarding the contribution of the surface, is then considered Gujrati ; MinosJurgen ; tiagoPol . Basically, one assumes that the free energy per site of sites located on the surface of the tree is $\phi_{s}$ and for sites in the bulk it is equal to $\phi_{b}$. The result for the bulk free energy per site is:
$\phi_{b}=-\frac{k_{B}T}{4}\ln\left[\frac{Y^{\prime}}{Y^{3}}\right],$ (16)
where $Y$ and $Y^{\prime}$ are the partition functions of Husimi trees with $N$ and $N+1$ generations, respectively. It may then be shown that
@@
In the calculations above we defined the Husimi tree so that the center of an elementary square is located at at its center. We could, alternatively, define it in such a way that a site of the tree is located at the center. In this case, although we would expect different results for a finite tree, the average values at the central region of both trees will be equal in the thermodynamic limit. The partition function in this case is obtained connecting two sub-trees with the same number of generations to the central site, giving
$Y_{a,N}=g_{0}^{2}+\sum_{n=0}^{k-1}[z_{x}g_{(n,x)}g_{(k-n-1,x)}+z_{y}g_{(n,y)}g_{(k-n-1,y)}]=g_{o}^{2}y_{a},$ (18)
@@
The average number of $k$-mers in the horizontal and vertical direction reaching the central site in the thermodynamic limit may be expressed in terms of the parameters $x_{1}$ and $x_{2}$ (Eq. 8) as
@@
and a procedure similar to the one used above for the square-centered tree leads to the bulk free energy per site:
@@
where $f_{0}$ is given in Eq. 12
@@
III Results for rods on the square HL

### v2

# Hard rigid rods on Husimi lattices

@@ 1 Introduction
Particle systems with excluded volume interactions only have been studied for much time. Besides a disordered phase, they may also exhibit ordered phases and thus undergo entropy-driven transitions between them. Many different shapes have been considered for the particles, such as hard spheres, squares, triangles, tetraminoes, and so on [1, 2, 3, 4, 5, 6, 7]. Some systems composed by binary [8] and ternary mixtures [9] of hard particles have also been considered in the literature. In this paper, we investigate hard rigid rods, which are frequently associated to liquid crystals [10]. The study of this particular system started with a seminal paper by Onsager [11], where he concluded that, at sufficiently high densities and large aspect ratios, the rods would undergo a discontinuous phase transition between an isotropic and a nematic phase. While Onsager considered the rods in continuous space, Flory [12] placed them on a lattice, but still allowing their orientations to be continuous. He solved the problem in a mean-field approximation, finding also a transition between isotropic and nematic phases. Zwanzig [13] studied a system of hard rods in continuum space, but with discrete orientations, finding, as Onsager, a discontinuous isotropic-nematic phase transition.
@@
Here, we are interested in rigid rods placed on lattices, formed by $k$ consecutive monomers aligned in one lattice direction, so that between consecutive monomers of a rod there is a lattice edge. We will call these rods $k$-mers. For the particular case of $k=2$ (dimers) on the square lattice, the entropy of the system in the limit of full coverage has been exactly determined, using both pfaffians [14, 15] and transfer matrices [16, 17], being $s_{2}^{(sq)}=G/\pi=0.29156\ldots$ the entropy per site, where $G$ is Catalan’s constant. Fully-packed dimers on the triangular lattice have also a long history (see, e.g., Refs. [18, 19, 20, 21, 22]), where $s_{2}^{(tri)}=0.42859\ldots$ [21, 22]. The exact determination of the entropy for dimers in this limit is also possible for other two-dimensional lattices [22], as well as in higher dimensions [23]. For larger values of $k$, as far as we known, there are no exact results available for the full lattice entropy of $k$-mers placed on regular lattices, except for the asymptotic $k\rightarrow\infty$ limit, where $s_{k}=k^{-2}\ln k$ on hypercubic lattices [41]. However, rather precise estimates have been obtained using transfer matrix methods for trimers [24] and other $k$-mers with $2\leq k\leq 10$ [25] on the square lattice, which agree with the outcomes from numerical simulations [26]. Semi-analytical solutions of fully-packed rods on generalized Husimi lattices — built with square lattice clusters of effective lateral size $L$ — also provide results in fair agreement with those for the square lattice, when extrapolated to the $L\rightarrow\infty$ limit [27]. To the best of our knowledgement, there are no analogous results for the triangular lattice reported in the literature.
@@
Turning our attention to the more general case where vacancies may exist between rods placed on the square or triangular lattices, we start mentioning that Monte Carlo simulations demonstrate that no nematic order exists for $k<k_{min}=7$, whereas a continuous isotropic-nematic transition is found for $k\geq 7$. Its universality class depends on the number of possible orientations of the rods; and may be Ising [29, 30, 31] or, more generally, Potts [29, 30, 32]. The existence of the nematic phase at intermediate densities was also rigorously established for large $k$’s [33]. Interestingly, a second transition — from the nematic to a high-density disordered phase — is found at high rods’ densities [28], but its nature is not so well known. Although usual Monte Carlo algorithms using local evaporation and deposition moves are very inefficient in the high density region, a new procedure introduced some time ago (simultaneously updating all sites of a strip using transfer matrix procedures) reduced this difficulty [34, 35], indicating a non-Ising transition for $k=7$. However, recent results suggest that this transition is discontinuous for large $k$ [36].
@@
The thermodynamic behavior of $k$-mers was also studied on the Bethe lattice (BL) and Bethe-like lattices [37]. There, a continuous isotropic-nematic transition was found for the four-coordinated lattice ($q=4$) already for $k\geq k_{min}^{(BL)}=4$; in contrast with $k_{min}=7$ obtained for the square lattice [28]. For BLs with coordination $q\geq 6$, the isotropic-nematic transition is discontinuous and appears, again, for $k\geq k_{min}^{(BL)}=4$; which differs from the continuous transition found (for $k\geq 7$) in the triangular lattice [30]. Furthermore, the transition from the nematic to the high density isotropic phase is missing in the BL for all rod size $k$ and coordination $q$. We recall that the exact (or numerically exact) solution of a given model on the BL corresponds to the result of the so-called mean-field Bethe approximation on a regular lattice with the same coordination number [38]. So, the differences above lead us to inquire whether the rods’ behavior can be better captured by more elaborated mean-field approaches. In order to address this, in this paper, we study the thermodynamic properties of $k$-mers, for general densities, on the four-coordinated Husimi lattice built with squares and on the six-coordinated Husimi lattice built with triangles. Since loops are present in these Husimi cacti, they are expected to provide better mean-field approximations for the models on the square and triangular lattices than the simpler BL solution [39, 40]. This was indeed the case in the full lattice limit analyzed in Ref. [27], as already mentioned above. As demonstrated in what follows, the scenario found here is not so different to that just described for the BL; even though, in most cases, we obtain critical parameters closer to those estimated in the literature for the regular lattices, provided by techniques such as Monte Carlo simulations or transfer matrix calculations.
@@
The rest of the paper is organized as follows. In Sec. 2 the Husimi lattices are defined and the model is solved on the one built with squares; while the solution on the triangular Husimi cactus is presented in the Appendix. The thermodynamic behavior of rods placed on the square and triangular Husimi lattices are shown and discussed in sections 3 and 4, respectively. Our final discussions may be found in Sec. 5.
2 Solution of the model on the Husimi lattice built with squares
The Husimi cactus with coordination number $q=4$, as consider here, is a hierarchical structure where successive generations (of shells) of squares are added to a central square, as is shown in Fig 1(a). A similar cactus, with coordination $q=6$, can be built up by joining generations of triangles to a central triangle, as done in Fig. 1(b). In contrast to regular lattices, the ratio between the number of surface and bulk sites does not vanish in these cacti even in the thermodynamic limit (i.e., when the number $N$ of generations diverges). However, the core of an infinity Husimi cactus — i.e., its inner region infinitely far away from the surface — is expected to form a lattice  [^1] See, e.g., Refs. [38, 52] for clear discussions on this in the context of the Bethe lattice, which is formed by connecting the centers of the polygons forming the Husimi lattices analized here., which is known in the literature as a Husimi lattice (HL) [40]. The solution of a given model in the square HL (triangular HL) illustrated in Fig. 1 provides a mean-field approximation for the model’s behavior on the square (triangular) lattice. Notice also that, since surface effects shall become irrelevant in these HLs, the model’s properties on them are expected to be shell-independent.
@@ 2 Solution of the model on the Husimi lattice built with squares
Figure 1: Husimi cactuses built with (a) squares and (b) triangles, by starting with a central polygon and adding two generations of them at the surface sites of the previous generations. Following this procedure, such cactuses always have an elementary polygon [square in (a) and triangle in (b)] at their centers. The four-coordinated Cayley tree obtained by connecting the squares’ centers in (a) is represented by dashed lines.
@@
With such considerations in mind, the solution of the rigid rods model on the square HL will be presented here, while the one for the triangular case will be discussed in appendix A. The model we are analyzing is a generalization of the full lattice limit, which was recently considered by some of us on the square HL [27], but now empty sites are allowed and activities $z_{x}=\exp[\mu_{x}/(k_{B}T)]$ and $z_{y}=\exp[\mu_{y}/(k_{B}T)]$ are associated to each site occupied by a monomer in a rod aligned in the $x$ and $y$ direction, respectively. The quantities $\mu_{x}$ and $\mu_{y}$ are the chemical potentials of a monomer, which is part of a $k$-mer, so that we are in the grand-canonical ensemble. Moreover, $k_{B}$ is Boltzmann’s constant and $T$ is the temperature. Although we will restrict ourselves to the thermodynamic properties of the system in the case $z_{x}=z_{y}=z$, for latter convenience, the solution will be presented considering different activities. As expected, in all expressions below the full lattice limit results are recovered in the $z\to\infty$ limit.
@@
As usual in solving models on such lattices, we start obtaining recursion relations for the grand-canonical partial partition functions (ppf’s) of rooted sub-cactuses for fixed configurations of their root site. In other words, we consider the operation of connecting three sub-cactuses to three corners of a new root square to build up a new sub-cactus with an additional generation of squares. We will label the root configuration of the sub-cactus according to the number of monomers already attached to the rod reaching it from above. If there is no such rod, this label will be equal to zero. If there is a rod reaching it, the label will be the number of monomers already included in the rod with a second component that indicates the direction of the rod, $x$ or $y$. Therefore, there are $12k-$ 1 ppf’s, whose first labels are in the range $[0,...,k-1]$. In Fig. 2 the possible sub-cactus root configurations are shown.
Figure 2: Possible root configurations of a sub-cactus. The thick red lines represent bonds of a $k$-mer on the corresponding edge, and the red dots are monomers of rods. Notice that the activity of the monomer which eventually is located at the root site will be considered in the next iteration and, for this reason, it is not represented in the figures. From the left to the right: no rod reaching on the root site from above ($g_{0}$), a rod in the $x$-direction arriving at the root site ($g_{(i,x)},\,i=1,2,...,k-1$), and a rod in the $y$-direction reaching on the root site ($g_{(i,y)},i=1,2,...,k-1$).
@@
We now proceed obtaining the recursion relations for the ppf’s, by considering the operation of attaching three sub-cactuses to a new root square. The possible configurations of the new root square for the case of no rod arriving at the new root site, are illustrated in Fig. 3.
@@
Figure 3: Possible configurations of the root square contributing to the recursion relation for $g^{\prime}_{0}$. The thick red lines indicate edges of a $k$-mer placed on the new root square, while its monomers are represented by the red dots.
@@
where it may be helpful to notice that when a bond belonging to a rod occupies one of the two edges in the upper part of the root square, the incoming two parts of this rod should have $n$ and $k-2-n$ monomers, respectively, with $n$ assuming values between 00 and $k-2$, which explains the sums in the two last terms of the recursion relation. The other recursion relations, for ppf’s of sub-cactuses with rods arriving at the root site, may be obtained using similar combinatorial arguments. The results are:
@@
for $i=1,2,...,k-1$. We notice that, if we write the recursion relations for the ppf’s as $g^{\prime}_{(i,x)}=g_{0}^{3}f_{(i,x)}$ and $g^{\prime}_{(i,y)}=g_{0}^{3}f_{(i,y)}$, the functions $f_{i}$ depend only of the activities and the ratios defined above. It is easy to obtain the $2(k-1)$ recursion relations for the ratios. Defining $R_{(i,x)}^{\prime}=g_{(i,x)}^{\prime}/{g_{0}}^{\prime}$ and $R_{(i,y}^{\prime}=g_{(i,y)}^{\prime}/g_{0}^{\prime}$, they are:
@@
The grand-canonical partition function $Y_{N}$ of the system may be obtained considering the operation of connecting four sub-cactuses with $N$ generations to the central square of the Husimi cactus. A convenient alternative is to connect a sub-cactus with $N+1$ generations to another one with $N$ generations, so that they share their root sites. This common root site may be empty or occupied by a monomer of a horizontal or vertical rod, thus:
@@
where we notice that the sum corresponds to the configurations where a monomer occupies the site shared by the both sub-cactuses, so that $n$ monomers of the rod to which this monomer belongs are on sites of the sub-cactus with $N+1$ generations and $k-n-1$ of its monomers are on sites of the other sub-cactus. The partition function is proportional to $g_{0}^{4}$, where it is implicit that this corresponds to a ppf of a sub-cactus with an empty root site and $N$ generations, and the proportionality factor is given by:
@@
The average number of $k$-mers in the horizontal and vertical directions reaching the site of the central square of the cactus where the two sub-cactuses meet are, respectively:
@@
The free energy of the system on the whole Husimi cactus, in the thermodynamic limit, is dominated by the surface. The bulk grand-canonical free energy per site, limited to the central region, discarding the contribution of the surface, is then considered [42, 43, 44]. Basically, one can assume that the free energy per site of sites located on the surface of the cactus is $\phi_{s}$ and for sites in the bulk it is equal to $\phi_{b}$. Then, one applies this assumption to two cactuses, with $N$ and $N+1$ generations respectively, which allows the determination of $\phi_{b}$ (as well as $\phi_{s}$). The result for the bulk free energy per site is:
$\phi_{b}=-\frac{k_{B}T}{4}\ln\left[\frac{Y_{N+1}}{Y_{N}^{3}}\right].$ (16)
We remark that the same result can be obtained, in the $N\rightarrow\infty$ limit, if one assumes that $\phi$ is a function of the cactus shell, being $\phi=\phi_{b}$ at the center, as above [51]. It may then be shown that
@@
In the calculations above we defined the Husimi cactus so that the center of an elementary square is located at its center. We could alternatively define it in such a way that a site of the cactus is located at the center. In this case, although we would expect different results for a finite cactus, the average values at the central region of both cactuses will be equal in the thermodynamic limit. The partition function in this case is obtained connecting two sub-cactuses with the same number of generations to the central site, giving
$Y_{a,N}=g_{0}^{2}+\sum_{n=0}^{k-1}[z_{x}g_{(n,x)}g_{(k-n-1,x)}+z_{y}g_{(n,y)}g_{(k-n-1,y)}]=g_{0}^{2}y_{a},$ (18)
@@
The average number of $k$-mers in the horizontal and vertical directions reaching the central site in the thermodynamic limit may be expressed in terms of the parameters $x_{1}$ and $x_{2}$ (Eq. 8) as
@@
and a procedure similar to the one used above for the square-centered cactus leads to the bulk free energy per site:
@@
where $f_{0}$ is given in Eq. 12.
One point which is worth looking at is whether the free energies per site [in Eqs. (17) and (21)] are consistent with the ones that would be obtained integrating the densities with respect to the chemical potential. So, considering that
$n_{x}=-k_{B}T\left(\frac{\partial\phi_{b}}{\partial\mu_{x}}\right)_{\mu_{y},T}=-z_{x}\left(\frac{\partial\phi_{b}}{\partial z_{x}}\right)_{z_{y}},$ (22)
and
$n_{y}=-z_{y}\left(\frac{\partial\phi_{b}}{\partial z_{y}}\right)_{z_{x}},$ (23)
we have calculated the densities of monomers in $k$-mers in both lattice directions directly from the free energy in Eq. (21) and compared the results with those provided by Eqs. (20) — for several values of the activity per monomer $z$ and of the number of monomers in rods $k$, both in the isotropic and in the nematic phase —, finding that they coincide within the numerical precision we are working.
@@
3 Results for rods on the square HL